# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v4

@@ H.2 Technical Lemma 2
For let $G_{n}(\zeta):=\|\zeta\|_{n}^{-1}\sum_{i=1}^{n}h\left(U_{i}\right)E_{i}$. Then by (47),
@@
Thus by (49) we have that
@@
Define the concave function $\Phi:=\widehat{\phi}-\widetilde{\phi}$. Now, note that by H.4 we have
@@
  • Step 1.A. Since $\Phi$ is concave, it follows that if $\Phi\left(t_{o}+\delta\right)\geq\xi/2$ then, joint with (51), $\Phi\geq\xi/2$ on $\left[t_{o},t_{o}+\delta\right]$.
@@
  • Step 1.A + 2.A: $\Phi\geq\xi/2$ or $\Phi\leq\xi/2$ on some interval $J\subset\left[t_{o},t_{o}+4\delta\right]$ with length $\delta$. Using this fact and (50), we have that $t\in J$
@@
    Now suppose that $(\phi-\widehat{\phi})\left(t_{o}\right)\geq\xi$ for some $\xi>0$ and $t_{o}\in[a+\delta,b-\delta]$, where $0<\delta\leq(b-a)/2$. By H.4 and concavity of both $\phi,\widehat{\phi}$ there exist numbers $\gamma,\widehat{\gamma}$ such that
@@ Appendix I Remarks
I.1 Remark on 3.2
In this section, we discuss sufficient conditions under which 3.2 holds. Under twice differentiability of $\psi_{i}$, the strong concavity of $\operatorname{rev}_{i}\left(\cdot\mid\mathbf{p}_{-i}\right)$ uniformly over $\mathbf{p}_{-i}$, is guaranteed if there exists a value $\mu_{i}>0$, such that $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})\geq\mu_{i}$ for every $p\in\mathcal{P}_{i}$, uniformly in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. Uniformity in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$ means that $\mu_{i}$ does not depends on $\mathbf{p}_{-i}$. Here we give some examples where this condition is satisfied.
@@
    Concave demand models. If $\psi_{i}$ is such that $\psi_{i}^{\prime}\geq B_{\psi_{i}^{\prime}}$ for some $B_{\psi_{i}^{\prime}}>0$ as in 3.1 and is concave, which implies $\partial_{p}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=\psi_{i}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})-\beta_{i}p\psi_{i}^{\prime}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, then
@@
    Note that requiring concavity of $\psi_{i}$ is equivalent to requiring 11-concavity of $\psi_{i}$, then, by the inclusion property of $s$-concave functions, any (positive and strictly increasing) $s_{i}$-concave $\psi_{i}$ with $s_{i}\geq 1$ satisfies 3.2.
@@
    s-concave demand functions. Suppose that $\psi_{i}$ satisfies 3.1 and 3.4. Then, by Proposition 3.5, $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, and by Lemma B.1 this is equivalent to $\psi_{i}^{\prime\prime}\leq(1-s_{i})(\psi_{i}^{\prime})^{2}/\psi_{i}$. This implies that
@@
I.2 Remark 4.1 on the common exploration phase
@@
I.3 Remark 4.2 on the need for two different phases for model estimation
@@
We evaluate the performance of Algorithm 1 in the sequential price competition model with $N\in\{2,4,6\}$ sellers. Each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$ sets prices supported on $\mathcal{P}=[0,3]^{\times N}$. We examine how convergence behavior changes as the contraction constant $L_{\bm{\Gamma}}$ varies –specifically, when it is close to 0, near $0.5$, or approaches $1$.
@@
which simplifies to $L_{\bm{\Gamma}}=\sup_{i\in\mathcal{N}}\frac{\left\|\gamma_{i}\right\|_{1}}{\beta_{i}}$, when the link functions $\psi_{i}$ are log-concave (see Remark 3.8). In our experiments, we take
@@
A summary of the two extreme cases can be found in Table 3.
@@
Given the discussion above, and the table in Table 3, for each $N\in\{2,4,6\}$, we run simulations under three representative values of $\beta_{i}$, yielding a total of $9$ settings:
@@
In all experiments, demand noise follows $\text{Unif}[-0.05,0.05]$. The common exploration phase has length $\frac{2}{3}T^{5/7}$ and the length $\kappa_{i}$ of the first part of the exploration phase is determined using Equation 14 with $\mathscr{C}_{i}=1$, $\mathscr{B}_{i}=1$. During the first phase of the exploration phase, the values of $\mathbf{p}^{(t)}$ are samples from a multivariate Gaussian with mean
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we apply Algorithm 1 and repeat the simulation $30$ times to obtain averages and $95\%$ confidence intervals.
The results are displayed in Figure 6, and a log-log version to highlight the convergence rate in Figure 6. As expected, convergence accelerates as the contraction constant approaches 11, in line with the intuition highlighted in (9).
@@
In this experiment, we replicate the setup in Section J.1 but relax the common exploration phase assumption in order to test the robustness of our algorithm. An illustration of misaligned exploration phases is provided in Figure 4.
Figure 4: Illustration of our policy (Algorithm 1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau_{i}$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box with dotted border), with length $\tau_{i}\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box with continued border line), with length $\tau_{i}(1-\kappa_{i})$. The $i$-th seller’s price experiment ends at period $t=\tau_{i}$ (black circles). Subsequently, in their exploitation phase (represented by the continued black line), seller $i$ offers prices based on the estimators generated in the exploration phase.
@@
where $\tau_{\text{base}}=\tfrac{2}{3}T^{5/7}$. The allocation parameter $\kappa_{i}$ for the first part of exploration is computed using Equation 14, with $\tau$ replaced by $\tau_{i}$ and fixed constants $\mathscr{C}_{i}=\mathscr{B}_{i}=1$. For this simulation, we set $\beta_{i}=0.8$ for $N=2$, $\beta_{i}=0.92$ for $N=4$, and $\beta_{i}=97$ for $N=6$, which satisfy 3.6.
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we run Algorithm 1 and repeat the simulation 30 times, reporting averages and $95\%$ confidence intervals.
The results are reported in Figure 7, together with a log-log plot in Figure 6, which illustrate the convergence of prices to the Nash equilibrium and the decay of regret.
As anticipated, the estimation of $\psi_{i}$ deteriorates, in line with the intuition of Remark 4.1. Indeed, as the horizon $T$ grows, the exploration lengths $\tau_{i}$ increase for all sellers, which enlarges the non-common exploration phase (see Figure 7). Consequently, a larger number of price vectors $\mathbf{p}^{(t)}$ are no longer i.i.d., making parameter estimation increasingly inconsistent. Nevertheless, despite the adverse effect on the estimation of $\psi_{i}$, both convergence to the Nash equilibrium and regret performance remain robust, with empirical regret rates matching – or even surpassing – the theoretical benchmarks of 1/7-1/7 for equilibrium convergence and $5/7$ for regret.
@@
In this experiment, we replicate the setup of Section J.1, except that we fix $N=4$ and set $\beta_{i}=0.9$. We again take $\psi_{i}(u)=\Phi(2u/i)$ for $i=1,2,\dots,N$, where $\Phi$ denotes the c.d.f. of a standard Gaussian random variable. Since $\Phi$ is log-concave, each $\psi_{i}$ is 0-concave, that is, $s_{i}=0$. We then run Algorithm 1 under five different shape parameters $s_{i}^{\prime}\in\{-0.2,-0.1,0,0.1,0.2\}$. As shown in Figure 8, this misspecification of $s_{i}$ has no visible effect on either the NE convergence rate or the regret convergence rate. The small variations in the estimated slopes reported in the legend of Figure 8 are plausibly due to finite-sample noise ($T=1600$).
@@
Figure 9: Illustration of our policy (Algorithm 1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box), with length $\tau\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box), with length $\tau(1-\kappa_{i})$. The seller’s price experiment ends at period $t=\tau$ (orange circle). Subsequently, in their exploitation phase (dashed line), seller $i$ offers prices using the estimators generated in the exploration phase.
@@
In bagnoli2006log we find a large class of log-concave CDF. They prove that if a density function $f$ is log-concave and continuously differentiable, then the CDF $F$ is also log-concave. This gives a way to generate log-concave CDFs. Figure 10, by bagnoli2006log, presents several widely used continuous univariate probability distributions whose densities are log-concave. Except for the Laplace distribution, the log-concavity of each density $f$ can be verified by checking that the second derivative of $\ln f(x)$ is non-positive across its support. For some families-such as the Weibull, power-function, beta, and gamma distributions-log-concavity holds only for specific parameter values. The table reports the parameter ranges under which these densities remain log-concave.
@@
Figure 11, by bagnoli2006log summarizes, for each of these cases, whether the density and the CDF are log-concave or log-convex.
@@
From Figure 11, we note that the CDFs of the Pareto, Lognormal. Student’s t, Cauchy are not log-concave. In the next section, we prove that they are $s^{*}$-optimal-concave for some $s^{*}<0$.
@@
In this section, we present some properties of $s$-concave functions and give some examples. In this Section, we will widely use Lemma B.1 without explicit mention, which states that if $f$ is a positive function defined in an interval $(a,b)$ that is twice continuously differentiable, then $f$ is $s$-concave iff
@@
From now on, we assume that the PDFs that we consider are continuous. As we showed in Section L.2, many density functions are not log-concave, such as Student’s t, Cauchy, Pareto, and log-Normal. However, for all of them, we can show that the PDFs are $s$-concave for some $s<0$.
@@
The proof of Proposition L.6 is deferred to Section L.3.5.
@@
The proof of Proposition L.7 is deferred to Section L.3.6.
@@
In the following proposition, we prove that, under certain conditions, when a density function is $s$-concave, then the CDF and the survival function is $\mu$-concave for some $\mu=\mu(s)$. The proof of Proposition L.8 is deferred to Section L.3.7.
@@
The proof of Proposition L.9 is deferred to Section L.3.8.
@@
From Figure 11 we already know that the CDFs of the Pareto, Lognormal. Student’s t, Cauchy are not log-concave. However, by our Proposition L.9 we immediately get that they are $s^{*}$-optimal-concave for some $s^{*}<0$.
@@
From Proposition L.8 (and Proposition L.9), if $f$ is $\mu^{\star}$-optimal-concave it does not necessarily means that $F$ (and $\bar{F}$) is $s(\mu^{*})=\left(1-\frac{1}{\mu^{*}+1}\right)$-optimal concave. Indeed, the optimal concave value can be larger than $s(\mu^{*})$. However, for the class of distributions in Corollary L.10, we know that $s^{*}$ has to be <0<0.

### v5

@@ H.2 Technical Lemma 2
For let $G_{n}(\zeta):=\|\zeta\|_{n}^{-1}\sum_{i=1}^{n}h\left(U_{i}\right)E_{i}$. Then by equation 47,
@@
Thus by equation 49 we have that
@@
Define the concave function $\Phi:=\widehat{\phi}-\widetilde{\phi}$. Now, note that by ˜H.4 we have
@@
  • Step 1.A. Since $\Phi$ is concave, it follows that if $\Phi\left(t_{o}+\delta\right)\geq\xi/2$ then, joint with equation 51, $\Phi\geq\xi/2$ on $\left[t_{o},t_{o}+\delta\right]$.
@@
  • Step 1.A + 2.A: $\Phi\geq\xi/2$ or $\Phi\leq\xi/2$ on some interval $J\subset\left[t_{o},t_{o}+4\delta\right]$ with length $\delta$. Using this fact and equation 50, we have that $t\in J$
@@
    Now suppose that $(\phi-\widehat{\phi})\left(t_{o}\right)\geq\xi$ for some $\xi>0$ and $t_{o}\in[a+\delta,b-\delta]$, where $0<\delta\leq(b-a)/2$. By ˜H.4 and concavity of both $\phi,\widehat{\phi}$ there exist numbers $\gamma,\widehat{\gamma}$ such that
@@ Appendix I Remarks
I.1 Remark on ˜3.2
In this section, we discuss sufficient conditions under which ˜3.2 holds. Under twice differentiability of $\psi_{i}$, the strong concavity of $\operatorname{rev}_{i}\left(\cdot\mid\mathbf{p}_{-i}\right)$ uniformly over $\mathbf{p}_{-i}$, is guaranteed if there exists a value $\mu_{i}>0$, such that $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})\geq\mu_{i}$ for every $p\in\mathcal{P}_{i}$, uniformly in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. Uniformity in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$ means that $\mu_{i}$ does not depends on $\mathbf{p}_{-i}$. Here we give some examples where this condition is satisfied.
@@
    Concave demand models. If $\psi_{i}$ is such that $\psi_{i}^{\prime}\geq B_{\psi_{i}^{\prime}}$ for some $B_{\psi_{i}^{\prime}}>0$ as in ˜3.1 and is concave, which implies $\partial_{p}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=\psi_{i}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})-\beta_{i}p\psi_{i}^{\prime}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, then
@@
    Note that requiring concavity of $\psi_{i}$ is equivalent to requiring 11-concavity of $\psi_{i}$, then, by the inclusion property of $s$-concave functions, any (positive and strictly increasing) $s_{i}$-concave $\psi_{i}$ with $s_{i}\geq 1$ satisfies ˜3.2.
@@
    s-concave demand functions. Suppose that $\psi_{i}$ satisfies ˜3.1 and ˜3.4. Then, by Proposition˜3.5, $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, and by Lemma˜B.1 this is equivalent to $\psi_{i}^{\prime\prime}\leq(1-s_{i})(\psi_{i}^{\prime})^{2}/\psi_{i}$. This implies that
@@
I.2 Remark˜4.1 on the common exploration phase
@@
I.3 Remark˜4.2 on the need for two different phases for model estimation
@@
We evaluate the performance of Algorithm˜1 in the sequential price competition model with $N\in\{2,4,6\}$ sellers. Each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$ sets prices supported on $\mathcal{P}=[0,3]^{\times N}$. We examine how convergence behavior changes as the contraction constant $L_{\bm{\Gamma}}$ varies –specifically, when it is close to 0, near $0.5$, or approaches $1$.
@@
which simplifies to $L_{\bm{\Gamma}}=\sup_{i\in\mathcal{N}}\frac{\left\|\gamma_{i}\right\|_{1}}{\beta_{i}}$, when the link functions $\psi_{i}$ are log-concave (see Remark˜3.8). In our experiments, we take
@@
A summary of the two extreme cases can be found in Table˜3.
@@
Given the discussion above, and the table in Table˜3, for each $N\in\{2,4,6\}$, we run simulations under three representative values of $\beta_{i}$, yielding a total of $9$ settings:
@@
In all experiments, demand noise follows $\text{Unif}[-0.05,0.05]$. The common exploration phase has length $\frac{2}{3}T^{5/7}$ and the length $\kappa_{i}$ of the first part of the exploration phase is determined using Equation˜14 with $\mathscr{C}_{i}=1$, $\mathscr{B}_{i}=1$. During the first phase of the exploration phase, the values of $\mathbf{p}^{(t)}$ are samples from a multivariate Gaussian with mean
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we apply Algorithm˜1 and repeat the simulation $30$ times to obtain averages and $95\%$ confidence intervals.
The results are displayed in Figure˜6, and a log-log version to highlight the convergence rate in Figure˜6. As expected, convergence accelerates as the contraction constant approaches 11, in line with the intuition highlighted in equation 9.
@@
In this experiment, we replicate the setup in Section˜J.1 but relax the common exploration phase assumption in order to test the robustness of our algorithm. An illustration of misaligned exploration phases is provided in Figure˜4.
Figure 4: Illustration of our policy (Algorithm˜1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau_{i}$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box with dotted border), with length $\tau_{i}\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box with continued border line), with length $\tau_{i}(1-\kappa_{i})$. The $i$-th seller’s price experiment ends at period $t=\tau_{i}$ (black circles). Subsequently, in their exploitation phase (represented by the continued black line), seller $i$ offers prices based on the estimators generated in the exploration phase.
@@
where $\tau_{\text{base}}=\tfrac{2}{3}T^{5/7}$. The allocation parameter $\kappa_{i}$ for the first part of exploration is computed using Equation˜14, with $\tau$ replaced by $\tau_{i}$ and fixed constants $\mathscr{C}_{i}=\mathscr{B}_{i}=1$. For this simulation, we set $\beta_{i}=0.8$ for $N=2$, $\beta_{i}=0.92$ for $N=4$, and $\beta_{i}=97$ for $N=6$, which satisfy ˜3.6.
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we run Algorithm˜1 and repeat the simulation 30 times, reporting averages and $95\%$ confidence intervals.
The results are reported in Figure˜7, together with a log-log plot in Figure˜6, which illustrate the convergence of prices to the Nash equilibrium and the decay of regret.
As anticipated, the estimation of $\psi_{i}$ deteriorates, in line with the intuition of Remark˜4.1. Indeed, as the horizon $T$ grows, the exploration lengths $\tau_{i}$ increase for all sellers, which enlarges the non-common exploration phase (see Figure˜7). Consequently, a larger number of price vectors $\mathbf{p}^{(t)}$ are no longer i.i.d., making parameter estimation increasingly inconsistent. Nevertheless, despite the adverse effect on the estimation of $\psi_{i}$, both convergence to the Nash equilibrium and regret performance remain robust, with empirical regret rates matching – or even surpassing – the theoretical benchmarks of 1/7-1/7 for equilibrium convergence and $5/7$ for regret.
@@
In this experiment, we replicate the setup of Section˜J.1, except that we fix $N=4$ and set $\beta_{i}=0.9$. We again take $\psi_{i}(u)=\Phi(2u/i)$ for $i=1,2,\dots,N$, where $\Phi$ denotes the c.d.f. of a standard Gaussian random variable. Since $\Phi$ is log-concave, each $\psi_{i}$ is 0-concave, that is, $s_{i}=0$. We then run Algorithm˜1 under five different shape parameters $s_{i}^{\prime}\in\{-0.2,-0.1,0,0.1,0.2\}$. As shown in Figure˜8, this misspecification of $s_{i}$ has no visible effect on either the NE convergence rate or the regret convergence rate. The small variations in the estimated slopes reported in the legend of Figure˜8 are plausibly due to finite-sample noise ($T=1600$).
@@
Figure 9: Illustration of our policy (Algorithm˜1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box), with length $\tau\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box), with length $\tau(1-\kappa_{i})$. The seller’s price experiment ends at period $t=\tau$ (orange circle). Subsequently, in their exploitation phase (dashed line), seller $i$ offers prices using the estimators generated in the exploration phase.
@@
In bagnoli2006log we find a large class of log-concave CDF. They prove that if a density function $f$ is log-concave and continuously differentiable, then the CDF $F$ is also log-concave. This gives a way to generate log-concave CDFs. Figure˜10, by bagnoli2006log, presents several widely used continuous univariate probability distributions whose densities are log-concave. Except for the Laplace distribution, the log-concavity of each density $f$ can be verified by checking that the second derivative of $\ln f(x)$ is non-positive across its support. For some families-such as the Weibull, power-function, beta, and gamma distributions-log-concavity holds only for specific parameter values. The table reports the parameter ranges under which these densities remain log-concave.
@@
Figure˜11, by bagnoli2006log summarizes, for each of these cases, whether the density and the CDF are log-concave or log-convex.
@@
From Figure˜11, we note that the CDFs of the Pareto, Lognormal. Student’s t, Cauchy are not log-concave. In the next section, we prove that they are $s^{*}$-optimal-concave for some $s^{*}<0$.
@@
In this section, we present some properties of $s$-concave functions and give some examples. In this Section, we will widely use Lemma˜B.1 without explicit mention, which states that if $f$ is a positive function defined in an interval $(a,b)$ that is twice continuously differentiable, then $f$ is $s$-concave iff
@@
From now on, we assume that the PDFs that we consider are continuous. As we showed in Section˜L.2, many density functions are not log-concave, such as Student’s t, Cauchy, Pareto, and log-Normal. However, for all of them, we can show that the PDFs are $s$-concave for some $s<0$.
@@
The proof of Proposition˜L.6 is deferred to Section˜L.3.5.
@@
The proof of Proposition˜L.7 is deferred to Section˜L.3.6.
@@
In the following proposition, we prove that, under certain conditions, when a density function is $s$-concave, then the CDF and the survival function is $\mu$-concave for some $\mu=\mu(s)$. The proof of Proposition˜L.8 is deferred to Section˜L.3.7.
@@
The proof of Proposition˜L.9 is deferred to Section˜L.3.8.
@@
From Figure˜11 we already know that the CDFs of the Pareto, Lognormal. Student’s t, Cauchy are not log-concave. However, by our Proposition˜L.9 we immediately get that they are $s^{*}$-optimal-concave for some $s^{*}<0$.
@@
From Proposition˜L.8 (and Proposition˜L.9), if $f$ is $\mu^{\star}$-optimal-concave it does not necessarily means that $F$ (and $\bar{F}$) is $s(\mu^{*})=\left(1-\frac{1}{\mu^{*}+1}\right)$-optimal concave. Indeed, the optimal concave value can be larger than $s(\mu^{*})$. However, for the class of distributions in Corollary˜L.10, we know that $s^{*}$ has to be <0<0.